Session: Nov 2001 (session 3)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Tonino Pisanu at Thu Nov 22 15:13:13 MET 2001) Success

Torun (Kaz Borkowski at Thu Nov 22 15:44:19 MET 2001) Minor failures
Till 9:08 recordings off source because of telescope control problems. The source fell below telescope horizon at 13:20.

Effelsberg (A. Kraus at Fri Nov 23 09:00:21 MET 2001) Minor failures
Due to a mistake of the operator, polarisations were swapped (only for this experiment). Between UT0655-0723 we stopped (strong wind).

Onsala (John Conway at Fri Nov 23 11:00:26 MET 2001) Minor failures
First half of expt missed due to high winds otherwise no known problems

Jodrell_Bank (Alastair Gunn at Sun Nov 25 20:58:28 MET 2001) Success
No known problems.